Suppose a market where demand is given by the equation Q = 30 - P, where Q is quantity and P is price in $. The offer is given at Q = 10 + 4P. Will the consumers, collectively considered, benefit from the authorities introducing a maximum price of $1 instead of having the price determined by a free market?

a) Yes, because consumers will pay a lower price than with a free market

b) Yes, because consumer surplus is increasing.

c) No, because state interference in the market always harms consumers.

d) No, because the consumer surplus is declining.
Answer #1

Demand curve is given by : Q= 30-P

Supply curve is given by : Q= 10+4P

Equilibrium price : Demand curve= Supply Curve

30-P=10+4P

20=5P ==> P=$4.

Since the maximum price(price ceiling) is $1, consumer will have to pay less amount than the equilibrium price($4) . So, the consumer will have to pay a lower price than with a free market so option 'A' is correct.

It must be noted that, price ceiling not necessarily always increase the consumer surplus.

Therefore , Option 'A' is correct.